# Sub-wavelength optical lattice in 2D materials

### v1

@@ Abstract
Recently, light-matter interaction has been vastly expanded as a control tool for inducing and enhancing many emergent non-equilibrium phenomena [1, 2]. However, conventional schemes for exploring such light-induced phenomena [3, 4, 4, 5, 6, 7, 8, 9, 10, 11, 12] rely on uniform and diffraction-limited free-space optics, which limits the spatial resolution and the efficiency of light-matter interaction. Here, we overcome these challenges using metasurface plasmon polaritons (MPPs) to form a sub-wavelength optical lattice. Specifically, we report a “nonlocal” pump-probe scheme where MPPs are excited to induce a spatially modulated AC Stark shift for excitons in a monolayer of MoSe2, several microns away from the illumination spot. Remarkably, we identify nearly two orders of magnitude reduction for the required modulation power compared to the free-space optical illumination counterpart. Moreover, we demonstrate a broadening of the excitons’ linewidth as a robust signature of MPP-induced periodic sub-diffraction modulation. Our results open new avenues for exploring power-efficient light-induced lattice phenomena below the diffraction limit in active chip-compatible MPP architectures.
.1 Main
Recent decades have witnessed a surge of interest in using light as a powerful tool to induce, control, and enhance exotic emergent non-equilibrium phenomena in quantum and topological materials [1, 2]. For example, light has been widely utilized for inducing AC Stark shift [3, 4, 5], Floquet bands [6, 7], signatures of magnetization [8, 9], superconductivity [10], anomalous Hall effect [11], and melting gapped state into semi-metallic ones [12]. Common schemes for exploring such light-induced phenomena are based on diffraction-limited free-space optical spectroscopic techniques, which do not provide spatial patterns to imprint sub-wavelength lattice features [13]. These methods fundamentally put a bound on the spatial resolution and the efficiency of light-matter interaction due to the limited confinement of free-space optical beams. Moreover, pump-probe spectroscopy of emergent light-induced phenomena typically is performed “locally”, i.e., the pump and the probe illuminate and investigate the same spot. Such local pump-probe techniques cause unavoidable parasitic thermal and pump-induced noise, which necessitates complex heat management and pump-rejection methods [14]. The development of novel platforms and spectroscopy techniques to address these challenges remains an active area of research [15].
@@
Here, we design and realize a novel MPP-based nonlocal pump-probe scheme for inducing sub-wavelength optical lattices. Specifically, the excited MPPs imprint a one-dimensional lattice for excitons, resulting from a spatially modulated AC Stark shift.
@@
Figure 1(c) shows the normalized electric field intensity $|E|^{2}$ in the x-z plane (shown only for one groove) at 800 nm obtained from finite-difference time-domain (FDTD) simulations. The electric field intensity is maximum at hot spots near the sharp corners of the air-silver interface and decays exponentially with distance from the surface; this is a characteristic feature of SPPs. The TMD is located $\sim 10$ nm above the surface, as marked by the dashed white line in Fig. 1(b). The red curve shows the electric field intensity $|E|^{2}$ at the monolayer’s position. Importantly, due to this sub-diffraction variation of $|E|^{2}$, the excitons experience a position-dependent AC stark shift: those located at the local maxima of $|E|^{2}$ experience a larger blue shift compared to other excitons (represented by the dark and light blue ellipses in Fig. 1(b); see SI for detailed study of spatial dependence of $|E|^{2}$ and the full scheme of the experimental setup).
The diffractionless propagation of MPPs is a fundamental advantage for our nonlocal AC Stark shift scheme. To understand this feature, we calculate the dispersion of the device using the FDTD method. Figure 1(d) shows the iso-frequency contours of the device for different wavelengths. Here, $k_{x}$ and $k_{y}$ are the wavevectors along the x and y directions. Notably, at wavelengths of $\sim 700-800$ nm, the MPPs propagate with maximally suppressed diffraction. Moreover, for lower wavelengths (blue lines) the dispersion is hyperbolic with a negative angle of refraction, whereas it is elliptical with a positive angle of refraction for longer wavelengths (red line). Our numerical simulations closely follow the analytical dispersion relationship of the MPP device given by the coupled waveguide formalism under a slowly-varying amplitude approximation [35, 36]: $k_{y}=k_{wg}+2t\cos(k_{x}a)$, where $k_{wg}$ is the propagation constant of the MPP waveguide lattice, and $t$ is the tunneling amplitude between nearest-neighbor grooves in the lattice (see SI for details).
@@
We begin by characterizing the transmission spectrum of the device before stacking the TMD on the top, by coupling a broadband laser into one of the couplers and measuring the output from the other. The results are shown in Fig. 2(b). The transmission is maximum for a wavelength of $\sim 800$ nm and it decreases as a function of the detuning, in agreement with the simulated iso-frequency contours (Fig. 1(d)). The inset shows the simulated normalized transmission spectrum for the device monitored after 3 $\mu$m propagation of the MPPs. At greater detuning from the peak at $\sim 800$ nm, the measured transmission is smaller than in the simulation, since at large diffraction angles the MPPs do not successfully reach the bottom coupler. The asymmetric nature of the normalized transmission can be attributed to the longer propagation lengths for the elliptical MPPs [28]. To experimentally characterize the diffraction in the MPP device, we measure the spatial pattern of the emission from the output coupler for three different wavelengths as shown in Fig. 2(c). The red line shows the position of the pump laser at the input coupler. We observe diffracted outputs at wavelengths of 770 nm and 828 nm, whereas there is minimal diffraction at 796 nm, which is very close to our designed pump wavelength. Simulation and details of this wavelength-dependent behavior are available in the SI. After the initial MPP device characterizations, we complete the integrated device by stacking the hBN-encapsulated MoSe2 monolayer on the MPP device. Figure 2(d) shows a false-color microscope image after the stacking procedure.
@@
To demonstrate the sub-diffraction nonlocal AC Stark shift scheme introduced in Fig. 1, we excite the MPP modes using a 150 fs pulsed laser at 800 nm wavelength, as shown schematically in Fig. 3(a). The pump excites the MPPs in several grooves that lie under our diffraction-limited pump (see the SI for a discussion on the multi-groove excitation). They travel along the y-direction of the device and modulate the MoSe2 monolayer located on top of the MPP device. Importantly, the excited MPPs have a spatial periodic intensity profile with sharp features below the diffraction limit (see SI for simulations of the mode profiles). To probe the AC stark shift induced by the MPPs, we perform absorption spectroscopy using a weak pulsed broadband laser placed 3 $\mu$m away from the pump in a reflection configuration. We note that the pump is red-detuned by $\sim 50$ nm from the excitonic energy to prevent photoluminescence.
@@
Next, we confirm the superior power efficiency of our nonlocal pump-probe scheme, which arises from MPP-induced strong field confinement, by considering the optical power required to generate the same AC Stark shift using the conventional local pump-probe counterpart. In the local case, the pump and the probe are spatially overlapped as shown in Fig. 3(d). We perform the local AC Stark shift measurement on a TMD placed on an unpatterned silver substrate (device D2) as shown in Fig. 3(e). We observe an identical AC shift of $\sim 0.65$ nm using $\sim 25$ mW of pump power. Remarkably, when compared to the MPP-induced nonlocal AC Stark shift with an effective pump power of 0.33 mW, the nonlocal AC Stark shift scheme requires $\sim 90$ times less pump power to observe a comparable AC Stark shift for the same detuning (see the SI for details). Numerically, we estimate that the maximum field enhancement at the position of the TMD for our nonlocal scheme can be as high as 130 times the diffraction-limited optical pump intensity (see SI for detailed calculation). The slight discrepancy between the experimental observation and numerical estimate is likely due to fabrication imperfections like surface roughness, rounded groove corners, and deviation of the TMD position from its nominal value in the z-direction. Moreover, this remarkable observed enhancement is further substantiated by measuring a similar nonlocal AC Stark shift effect on a different device, which highlights the robustness of our scheme (see the data for device D4 in the SI). Furthermore, we confirm that the MPP-enabled strong field confinement is central to our scheme compared to 2D sheets (discussed earlier in Fig. 1(e-g)). To achieve this, we perform the same nonlocal scheme on a similar TMD device placed on a 2D sheet. Using an identical pump power strength, no AC Stark shift is observed in this case (see the nonlocal measurement of device D2 in the SI).
@@
We demonstrate another intriguing optical lattice-induced effect, in the form of an excitonic linewidth broadening, presented in Fig. 3(c). The excitons under our diffraction-limited probe laser experience a strong spatial variation of the electric field intensity of the MPPs as shown in Fig. 3(a). This periodic sub-diffraction modulation of electric field intensity (which can vary as strongly as $\sim 50$% as shown in the inset of Fig. 3(a)) manifests as a linewidth broadening of $\sim 1$ nm at $\Delta t=0$. This observation is in sharp contrast with the local counterpart, where there is no noticeable change in the linewidth at $\Delta t=0$ (Fig. 3(f)) (for a similar observation and discussion for a TMD monolayer placed on an unpatterned silica substrate, see the data for the device D3 in the SI).
We further confirm that this linewidth broadening originates from the sub-diffraction spatial modulation of the MPP by performing local AC Stark shift measurements on the top of the MPP device D1 (see Fig. S17). We observe a comparable linewidth broadening to the nonlocal case arising from the sub-diffraction modulation provided by the device, albeit, using a much larger pump power than the nonlocal case. Moreover, we note that thermal effects, which take place in much slower timescales than our short 150 fs pulse duration, are not likely to play a significant role in this observation. For the pump powers and detunings used in our experiments, we can also rule out possible contributions from power broadening (see the SI for detailed analysis).
@@
In summary, we demonstrated a novel TMD-integrated plasmonic metasurface approach for efficient modulation of excitons below the diffraction limit. Due to its intrinsic nonlocal pump-probe character [37], our platform provides a useful protocol for spatial on-chip pump rejection as well as suppressing parasitic thermal effects [14]. More generally, our results suggest that lattice SPPs in MPP device architectures can become a new and versatile toolbox to actively manipulate matter excitations below the diffraction limit.
Beyond our exemplary AC Stark shift case, a wide range of other light-induced phenomena can be investigated below the diffraction limit with superior power efficiency. This includes light-induced magnetization [8], superconductivity [38], quantum Hall effects [11], electronic polarization [39], polariton modulators [40] and plasmonic effects in van der Waals heterostructures [41, 42]. Another intriguing direction is the sub-wavelength periodic modulation of materials and matter excitations harnessing the versatile tunability and controllability of our MPPs. An example includes optical imprinting of lattices [13], where the optical drive forms a lattice structure in space. The lattice nature of MPPs also may enable engineering chiral light-matter interaction [31] using plasmonic spin Hall states [28].
In addition to light-induced phenomena, our platform could be useful for cavity-induced physics [2, 43, 44, 18]. In particular, the strong spatial confinement of light in such plasmonic structures combined with the hyperbolic photonic modes of the lattice can lead to intriguing phenomena such as meditated long-range interactions between distant emitters [45, 46].
@@
The authors wish to acknowledge fruitful discussions with Aaron Sternbach, Zubin Jacob, Andrey Grankin, Deric Session, and Nick Martin. This research was supported by DARPA HR00112490310, the National Science Foundation DMR-2145712, and the Department of Energy DE-SC-0022885 grant. This research used Quantum Material Press (QPress) of the Center for Functional Nanomaterials (CFN), which is a US Department of Energy, Office of Science User Facility, at Brookhaven National Laboratory under contract no. DE-SC0012704. K.W. and T.T. acknowledge support from the JSPS KAKENHI (grant nos. 20H00354, 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan, for hBN synthesis.
.8 Competing interests

### v2

@@ Abstract
Recently, light-matter interaction has been vastly expanded as a control tool for inducing and enhancing many emergent non-equilibrium phenomena. However, conventional schemes for exploring such light-induced phenomena rely on uniform and diffraction-limited free-space optics, which limits the spatial resolution and the efficiency of light-matter interaction. Here, we overcome these challenges using metasurface plasmon polaritons (MPPs) to form a sub-wavelength optical lattice. Specifically, we report a “non-local” pump-probe scheme where MPPs are excited to induce a spatially modulated AC Stark shift for excitons in a monolayer of MoSe2, several microns away from the illumination spot. Remarkably, we identify nearly two orders of magnitude reduction for the required modulation power compared to the free-space optical illumination counterpart. Moreover, we demonstrate a broadening of the excitons’ linewidth as a robust signature of MPP-induced periodic sub-diffraction modulation. Our results will allow exploring power-efficient light-induced lattice phenomena below the diffraction limit in active chip-compatible MPP architectures.
.1 Introduction
Recent decades have witnessed a surge of interest in using light as a powerful tool to induce, control, and enhance exotic emergent non-equilibrium phenomena in quantum and topological materials [1, 2]. For example, light has been widely utilized for inducing AC Stark shift [3, 4, 5, 6, 7, 8], Floquet bands [9, 10], signatures of magnetization [11, 12], superconductivity [13], anomalous Hall effect [14], and melting gapped state into semi-metallic ones [15]. Common schemes for exploring such light-induced phenomena are based on diffraction-limited free-space optical spectroscopic techniques, which do not provide spatial patterns to imprint sub-wavelength lattice features [16]. These methods fundamentally put a bound on the spatial resolution and the efficiency of light-matter interaction due to the limited confinement of free-space optical beams. Moreover, pump-probe spectroscopy of emergent light-induced phenomena typically is performed “locally”, i.e., the pump and the probe illuminate and investigate the same spot. Such local pump-probe techniques cause unavoidable parasitic thermal and pump-induced noise, which necessitates complex heat management and pump-rejection methods [17]. The development of innovative platforms and spectroscopy techniques to address these challenges remains an active area of research [18].
@@
Here, we design and realize an MPP-based non-local pump-probe scheme for inducing sub-wavelength optical lattices. Specifically, the excited MPPs imprint a one-dimensional lattice for excitons, resulting from a spatially modulated AC Stark shift.
@@
Figure 1(c) shows the normalized electric field intensity $|E|^{2}$ in the x-z plane (shown only for one groove) at 800 nm obtained from finite-difference time-domain (FDTD) simulations. The electric field intensity is maximum at hot spots near the sharp corners of the air-silver interface and decays exponentially with distance from the surface; this is a characteristic feature of SPPs. The TMD is located $\sim 10$ nm above the surface, as marked by the dashed white line in Fig. 1(b). The red curve shows the electric field intensity $|E|^{2}$ at the monolayer’s position. Importantly, due to this sub-diffraction variation of $|E|^{2}$, the excitons experience a position-dependent AC stark shift: those located at the local maxima of $|E|^{2}$ experience a larger blue shift compared to other excitons (represented by the dark and light blue ellipses in Fig. 1(b); see Supplementary Information (SI) for detailed study of spatial dependence of $|E|^{2}$ and the full scheme of the experimental setup).
The diffractionless propagation of MPPs is a fundamental advantage for our non-local AC Stark shift scheme. To understand this feature, we calculate the dispersion of the device using the FDTD method. Figure 1(d) shows the iso-frequency contours of the device for different wavelengths. Here, $k_{x}$ and $k_{y}$ are the wavevectors along the x and y directions. Notably, at wavelengths of $\sim 700-800$ nm, the MPPs propagate with maximally suppressed diffraction. Moreover, for lower wavelengths (blue lines) the dispersion is anisotropic and hyperbolic with a negative angle of refraction, whereas it is elliptical with a positive angle of refraction for longer wavelengths (red line). Our numerical simulations closely follow the analytical dispersion relationship of the MPP device given by the coupled waveguide formalism under a slowly-varying amplitude approximation [38, 39]: $k_{y}=k_{wg}+2t\cos(k_{x}a)$, where $k_{wg}$ is the propagation constant of the MPP waveguide lattice, and $t$ is the tunneling amplitude between nearest-neighbor grooves in the lattice (see SI for details). This behavior is analogous to the recent observation of canalized polaritons originating from flat iso-frequency contours in twisted MoO3 heterostructures [40, 41].
@@
We begin by characterizing the transmission spectrum of the device before stacking the TMD on the top, by coupling a broadband laser into one of the couplers and measuring the output from the other. The results are shown in Fig. 2(b). The transmission is maximum for a wavelength of $\sim 800$ nm and it decreases as a function of the detuning, in agreement with the simulated iso-frequency contours (Fig. 1(d)). The inset shows the simulated normalized transmission spectrum for the device monitored after 3 $\mu$m propagation of the MPPs. At greater detuning from the peak at $\sim 800$ nm, the measured transmission is smaller than in the simulation, since at large diffraction angles the MPPs do not successfully reach the bottom coupler. The asymmetric nature of the normalized transmission can be attributed to the longer propagation lengths for the elliptical MPPs [31]. We further note the observation of oscillations in the transmission spectra, possibly originating from the confinement of light by the ends of the 20 $\mu$m long MPP device forming a Fabry-Perot mode. To experimentally characterize the diffraction in the MPP device, we measure the spatial pattern of the emission from the output coupler for three different wavelengths as shown in Fig. 2(c). The red line shows the position of the pump laser at the input coupler. We observe diffracted outputs at wavelengths of 770 nm and 828 nm, whereas there is minimal diffraction at 796 nm, which is very close to our designed pump wavelength. Simulation and details of this wavelength-dependent behavior are available in the SI. After the initial MPP device characterizations, we complete the integrated device by stacking the hBN-encapsulated MoSe2 monolayer on the MPP device. Figure 2(d) shows a false-color microscope image after the stacking procedure.
@@
To demonstrate the sub-diffraction non-local AC Stark shift scheme introduced in Fig. 1, we excite the MPP modes using a 150 fs pulsed laser at 800 nm wavelength, as shown schematically in Fig. 3(a). The pump excites the MPPs in several grooves that lie under our diffraction-limited pump (see the SI for a discussion on the multi-groove excitation). They travel along the y-direction of the device and modulate the MoSe2 monolayer located on top of the MPP device. Importantly, the excited MPPs have a spatial periodic intensity profile with sharp features below the diffraction limit (see SI for simulations of the mode profiles). To probe the AC stark shift induced by the MPPs, we perform absorption spectroscopy using a weak pulsed broadband laser placed 3 $\mu$m away from the pump in a reflection configuration. We note that the pump is red-detuned by $\sim 50$ nm from the excitonic energy to prevent photoluminescence.
@@
Next, we confirm the superior power efficiency of our non-local pump-probe scheme, which arises from MPP-induced strong field confinement, by considering the optical power required to generate the same AC Stark shift using the conventional local pump-probe counterpart. In the local case, the pump and the probe are spatially overlapped as shown in Fig. 3(d). We perform the local AC Stark shift measurement on a TMD placed on an unpatterned silver substrate (device D2) as shown in Fig. 3(e). We observe an identical AC shift of $\sim 0.65$ nm using $\sim 25$ mW of pump power. Remarkably, when compared to the MPP-induced non-local AC Stark shift with an effective pump power of 0.33 mW, the non-local AC Stark shift scheme requires $\sim 90$ times less pump power to observe a comparable AC Stark shift for the same detuning (see SI for details). Numerically, we estimate that the maximum field enhancement at the position of the TMD for our non-local scheme can be as high as 130 times the diffraction-limited optical pump intensity (see SI for detailed calculation). The slight discrepancy between the experimental observation and numerical estimate is likely due to fabrication imperfections like surface roughness, rounded groove corners, and deviation of the TMD position from its nominal value in the z-direction. Moreover, this remarkable observed enhancement is further substantiated by measuring a similar non-local AC Stark shift effect on a different device, which highlights the robustness of our scheme (see the data for device D4 in SI). Furthermore, we confirm that the MPP-enabled strong field confinement is central to our scheme compared to 2D sheets (discussed earlier in Fig. 1(e-g)). To achieve this, we perform the same non-local scheme on a similar TMD device placed on a 2D sheet. Using an identical pump power strength, no AC Stark shift is observed in this case (see the non-local measurement of device D2 in SI).
@@
We demonstrate another intriguing optical lattice-induced effect, in the form of an excitonic linewidth broadening, presented in Fig. 3(c). The excitons under our diffraction-limited probe laser experience a strong spatial variation of the electric field intensity of the MPPs as shown in Fig. 3(a). This periodic sub-diffraction modulation of electric field intensity (which can vary as strongly as $\sim 50$% as shown in the inset of Fig. 3(a)) maps to a periodic modification of AC Stark shift of excitons. Excitons located at the maximum of the electric field intensity experience greater blue shift compared to excitons in their vicinity. This leads to inhomogeneous AC Stark shift of the excitons, thereby resulting in a linewidth broadening of $\sim 1$ nm at $\Delta t=0$. A theoretical model for the expected linewidth broadening resulting from the spatial profile of the electric field intensity is provided in SI and closely matches our experimental observations. This observation is in sharp contrast with the local counterpart, where there is no noticeable change in the linewidth at $\Delta t=0$ (Fig. 3(f)) (for a similar observation and discussion for a TMD monolayer placed on an unpatterned silica substrate, see the data for the device D3 in SI).
@@
In summary, we demonstrated a TMD-integrated plasmonic metasurface approach for efficient modulation of excitons below the diffraction limit. Due to its intrinsic non-local pump-probe character [42], our platform provides a useful protocol for spatial on-chip pump rejection as well as suppressing parasitic thermal effects [17]. More generally, our results suggest that lattice SPPs in MPP device architectures can become an innovative and versatile toolbox to actively manipulate matter excitations below the diffraction limit.
Beyond our exemplary AC Stark shift case, a wide range of other light-induced phenomena can be investigated below the diffraction limit with superior power efficiency and pump noise suppression. This includes light-induced magnetization [11], superconductivity [43], quantum Hall effects [14], electronic polarization [44], polariton modulators [45] and plasmonic effects in van der Waals heterostructures [46, 47]. Another intriguing direction is the sub-wavelength periodic modulation of materials and matter excitations harnessing the on-demand spectral and spatial tunability of our MPPs. An example includes optical imprinting of lattices [16], where the optical drive forms a lattice structure in space. The lattice nature of MPPs also may enable engineering chiral light-matter interaction [34] using plasmonic spin Hall states [31]. Moreover, the sub-diffraction spatial modulation and the electric field confinement can be probed using near-field optical microscopy for a better understanding of the MPP properties, and allow further innovative integration with solid-state emitters [48, 49, 50, 18].
In addition to light-induced phenomena, our platform could be useful for cavity-induced physics [2, 51, 52, 21]. In particular, the strong spatial confinement of light in such plasmonic structures combined with the hyperbolic photonic modes of the lattice can lead to intriguing phenomena such as meditated long-range interactions between distant emitters [53, 54, 55].
@@
The authors wish to acknowledge fruitful discussions with Aaron Sternbach, Zubin Jacob, Andrey Grankin, Deric Session, and Nick Martin.
.8 Funding
S.S., M.J.M., D.G.S.F., C.F., L.X., M.H. were supported by ARO W911NF2510066, DARPA HR00112490310 and HR00112530313. L.G. and Y.Z. were supported by ARO W911NF2510066, NSF DMR-2145712, and the Department of Energy (DOE) DE-SC-0022885 grants. This research used Quantum Material Press (QPress) of the Center for Functional Nanomaterials (CFN), which is a US Department of Energy, Office of Science User Facility, at Brookhaven National Laboratory under contract no. DE-SC0012704. K.W. and T.T. acknowledge support from the JSPS KAKENHI (grant nos. 20H00354, 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan, for hBN synthesis.
.9 Competing interests